# JailbreakLens: Interpreting Jailbreak Mechanism in the Lens of Representation and Circuit

### v1

@@ 5.4.1 Key Components Location
Figure 5: (a). Average refusal score attribution for each MLP layer in Llama2-7b on harmful and safe prompts. The $22_{nd}$ MLP layer shows a significant contribution to affirmations on safe prompts, while no layer demonstrates a significant contribution to refusals on harmful prompts. (b). Average refusal score attribution for each MLP layer in Llama2-7b on jailbreak prompts with the line of each color representing a different jailbreak method.
As shown in Fig. 4(a), only a small number of attention heads have a significant impact on safety in Llama2-7b. Specifically, the $14_{th}$ attention head in the $21_{st}$ layer (L21H14) in Llama2-7b has a strong contribution to refusing harmful instructions which we denote as the refusal signal head (i.e., $S_{-}$). While the $4_{th}$ head in the $26_{th}$ layer (L26H04) seems to enhance the affirmation response, which we term the affirmation signal head (i.e., $S_{+}$). Moreover, as shown in Fig. 5(a), the $22_{nd}$ MLP in Llama2-7b plays the most critical role in producing affirmative responses (i.e., the reverse of refusal), and the contribution of individual MLP layers to the refusal score on harmful prompts is relatively minor. Early studies [41, 42] on the importance of transformers components highlighted the model’s sparsity, meaning that removing a substantial fraction of attention heads may not degrade model performance. Our findings reinforce the previous observations, confirming that only a small subset of model components are crucial to generation safety.
Observation 5. Only a small number of model components have significant impacts on safety, with L21H14 as the refusal signal head, L26H04 as the affirmation signal head, and L22 MLP layer as the safety signal MLP layer. Moreover, the key components with a high contribution to safety are mainly located in the later layers.
Explaining the behaviors of key components in vocabulary space. To understand the specific role of each key attention head we identified, we map the activations on the signal heads of both safe prompts and harmful prompts to the vocabulary space, where the results on Llama2-7b are shown in Tab. III.
For the activations of refusal signal head $S_{-}$ on harmful prompts, the top decoded tokens include ‘ap’, ‘sorry’, and ‘exc’, indicative of apologetic or rejecting language. On safe prompts, $S_{+}$ activates tokens like ‘Hi’, ‘Cong’, and ‘welcome’, which reflect more neutral or polite expressions, but lack strong affirmative signals. In contrast, the affirmation signal head $S_{+}$ produces tokens like ‘tags’, ‘unas’, and ‘uz’ on harmful prompts, which seem scattered or unrelated. However, on safe prompts, $S_{+}$ returns tokens like ‘Abstract’, ‘Introduction’, and ‘Notice’, affirming the content in a more structured and informative way. These results indicate that $S_{-}$ effectively captures refusal-related language, while $S_{+}$ aligns more with affirmative language, demonstrating their distinct roles in identifying refusal and affirmation in the context of safe and harmful prompts.
Observation 6. Refusal signal heads effectively identify apologetic or rejecting tokens in response to harmful prompts, while affirmation signal heads capture guiding and informative tokens for safe prompts, highlighting their distinct roles in safeguarding generated outputs.
@@
TABLE III: Decoding the activation of signal attention heads in Llama2-7b for harmful and safe prompts in the vocabulary space. The activations of $S_{-}$ on harmful prompts primarily correspond to apologetic or rejecting tokens, while the activations of $S_{+}$ on safe prompts are largely associated with guiding or informative tokens.
Prompt Method Llama2-7b Vicuna1.5-7b
$S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$
Baseline Safe 1.0 -0.0760 1.0 0.4343
Harmful -0.0241 1.0 0.1562 1.0
Jailbreak GCG 0.9819 1.0 0.6021 1.0
ReNellm 1.0 0.3816 0.7485 1.0
PAIR 0.9810 1.0 0.5170 0.6897
DeepInception 1.0 -0.0997 1.0 0.7607
Autodan 0.0202 0.7573 0.4675 0.6523
GPTFuzz 0.0702 0.2817 0.4643 1.0
CodeChameleon 1.0 0.4816 0.4983 1.06
TABLE IV: Refusal attribution of signal attention head of Llama-2, and Vicuna-1.5 models with 7b scale on different jailbreak methods.
(a) Evolution of key components activation on a safe prompt.
(b) Evolution of representation on a safe prompt.
(c) Evolution of key components activation on a harmful prompt.
(d) Evolution of representation on a harmful prompt.
(e) Evolution of key components activation on DeepInc prompt.
(f) Evolution of representation on DeepIncprompt.
(g) Evolution of key components activation on Autodan prompt.
(h) Evolution of representation on Autodan prompt.
Figure 6: Dynamics of key circuit activation and representation on safe prompt, direct harmful prompt, and jailbreak prompts.
5.4.2 Effect of Jailbreak Prompts on Key Components
@@
After locating the key components that have a significant contribution to model safeguard, we then investigate how these components are activated when meeting jailbreak prompts. For ease of quantification, we normalize the activation values of each attention head to a range of $[-1,1]$ by dividing all the activation values by the maximum value among them, where positive values indicate enhancement of the signal, while negative values indicate suppression. The results of key attention heads and MLP layers are shown in Tab. IV and Fig. 5(b), respectively.
@@
For the signal attention heads, as summarized in Tab. IV, compared to their behavior on harmful prompts, all jailbreak methods suppress the activation of the refusal signal and enhance the activation of the affirmation signal. Specifically, for demonstration-based and rule-based methods, e.g., DeepInception, the refusal signal component is completely suppressed and the affirmation signal component is strongly enhanced, i.e., with $S_{+}$ enhancing from -0.0241 to 1.0 and $S_{-}$ suppressing from 1.0 to -0.0997. Evolution-based methods, such as GPTFuzz, also show significant suppression of the refusal signal, with $S_{-}$ decreasing from 1.0 to 0.2817, accompanied by a tiny enhancement of the affirmation signal, with $S_{+}$ increasing from -0.0241 to 0.0702. For gradient-based and multi-agent-based jailbreak methods, such as GCG, the suppression of refusal signals is less pronounced compared to harmful prompts, with no decrease of $S_{-}$, but the enhancement of the affirmation signal remains significant, with $S_{+}$ increasing from -0.0241 to 0.9819.
@@
Moreover, for the MLP layers, as illustrated in Fig. 5(b), most methods show similar behavior to the safe prompts, i.e., with suppression of the key MLP layer, the $22_{th}$ MLP. ReNellm demonstrates the strongest suppression of refusal on this MLP layer, followed by DeepInception, CodeChameleon, and GCG. In contrast, Autodan and GPTFuzz still show refusal enhancement on this MLP layer.
@@
Observation 7. Compared with harmful prompts, all the jailbreak strategies suppress the refusal signal components and enhance the affirmation signal components, while different strategy has different impact levels.
5.4.3 Impact of Fine-tuning on Key Components
As is well-known, the Vicuna series models are derived from instruction tuning applied to the Llama series. By comparing the results between the Llama and Vicuna series, we can assess the effect of fine-tuning, specifically instruction tuning, on the jailbreak mechanism. Notably, as shown in Fig. 4, both the 7b-scale Llama and Vicuna models exhibit the same locations of key attention heads, i.e., with L21H14 contributing significantly to the refusal function and L26H04 to the affirmation function.
@@ 5.4.3 Impact of Fine-tuning on Key Components
Furthermore, as shown in Fig. 7 in Appendix. B.1, the results for the 13b-scale models are consistent, with L37H37 and L31H35 serving as the locations for affirmation and refusal signal attention heads in both Llama2-13b and Vicuna1.5-13b. This consistency suggests that instruction tuning has a minimal impact on the functional components of the model’s circuit related to the refusal mechanism.
Observation 8. Instruction tuning has minimal impact on the model’s core refusal circuit, as the key attention heads linked to refusal and affirmation functions remain consistent across both Llama and Vicuna series models, indicating limited fine-tuning effects on internal response mechanisms.
5.5 Results of Interpretation Aggregation
In this section, we analyze interpretation results from representation and circuit analysis on the first token. Then we explore the evolution of key circuits and representations during the entire response generation process. For the representation analysis, we record the probing results of the later-layer representation (specifically, the $25_{th}$ layer) for each token generated by the model. Regarding the circuits, we track the activation of the signal attention heads identified in the previous section during the generation of each token in the model’s response. For clarity, partial results on Llama2-7b are presented here, with more detailed experimental results provided in Appendix  B.2.
@@ 5.5 Results of Interpretation Aggregation
Interpretation Aggregation on first token. Combining the analysis at the representation level, as shown in Fig. 2(b), we observe that jailbreak methods with a stronger impact on key components, specifically, those that more effectively suppress refusal signals and enhance affirmation signals, typically exhibit greater deception at the representation level.
@@
For example, this is evident in the comparison between DeepInception and GPTFuzz in Llama2-7b. As shown in Tab. IV, the activation of $S_{+}$ on DeepInception prompts is 1.0, the same as that on safe prompts, while the activation of $S_{-}$ is -0.0997, even lower than that on safe prompts. In contrast, the activation of $S_{+}$ on GPTFuzz prompts is 0.0702, indicating that the enhancement of the affirmation signal in GPTFuzz is significantly lower than in DeepInception. Meanwhile, the activation of $S_{-}$ for GPTFuzz is 0.2817, suggesting that the suppression of refusal signals in GPTFuzz is considerably higher than in DeepInception. Furthermore, as illustrated in Fig. 2(b), DeepInception prompts are predicted as safe by the probe in the final layers with high probability, while GPTFuzz prompts lie on the boundary between safe and harmful, indicating that the deception of DeepInception substantially stronger than that of GPTFuzz at the representation level.
@@
These findings suggest that methods like DeepInception, which effectively manipulate the attention mechanisms to suppress negative signals while reinforcing positive ones, create representations that align closely with the desired outputs of safe prompts. In contrast, GPTFuzz’s less effective manipulation leads to representations that do not fully disguise their harmful nature. This alignment between representation-level behavior and component functionality ultimately fosters a more comprehensive understanding of the interplay between representation and model components.
Observation 9. Jailbreak methods with a stronger impact on key components typically exhibit greater deception at the representation level. Although jailbreak shifts model representations toward safe clusters to deceive the LLM, leading it to provide detailed responses instead of refusals, it still produce abnormal activation which can be caught in the circuit analysis.
@@
Interpretation Aggregation thorough the entire generation process. We first evaluate the dynamic evolution of both safe and direct harmful prompts. Fig. 6(a) and Fig. 6(b) illustrate changes in signal head activation and representation throughout the generation process for a safe prompt on Llama2-7b. As shown in Fig. 6(a), in response to safe prompts, the affirmation signal remains consistently active, while the refusal signal is suppressed throughout the entire generation. Likewise, as shown in Fig. 6(b), the representations remain affirmative across the entire token sequence. The response dynamics for a harmful prompt are depicted in Fig. 6(c) and Fig.  6(d). In Fig.  6(c), the refusal signal is significantly amplified in the early stages of the response token sequence to the harmful prompt, while the affirmation signal is noticeably suppressed during the same period. As shown in Fig.  6(d), the representations of the first few tokens are refusal-oriented, with the model typically generating content such as “Sorry, I cannot fulfill your request…”. This activation pattern clearly demonstrates the model’s ability to reliably maintain a non-harmful response when interacting with malicious instructions.
@@
Fig. 10(k) and Fig. 10(l) provide an example of an Autodan prompt that fails to deceive the model at the representation level, leading to the successful generation of a refusal token as the initial token. Initially, the refusal signal is clearly enhanced, as shown in Fig. 10(k), and the representations of generated tokens take on negative values in the affirmation direction, as depicted in  10(l). The first few words generated by the model are typically refusal content, i.e., “Sorry, I cannot …”, which is similar to the response to the harmful prompt. However, after a certain point, the refusal signal is no longer amplified, and the representations stop exhibiting negative values in the affirmation direction. Based on the specific content generated by the model, it begins producing useful instruction on the harmful task required in the jailbreak prompt at this stage.
@@
Observation 10. In highly deceptive jailbreak prompts, the refusal signal is completely suppressed, resulting in the direct generation of harmful content. Conversely, an initial refusal in the response does not guarantee safety. In less deceptive jailbreak prompts, while the refusal signal may be activated at the beginning, it can either be bypassed immediately or gradually weakened throughout the entire generation process.
6 Discussion
Ethical Consideration. Our work aims to enhance the transparency and interpretability of LLMs by providing a detailed understanding of jailbreak mechanisms. By dissecting how jailbreak prompts disrupt safety measures and influence model behavior, we seek to identify vulnerabilities within LLMs that could lead to harmful outputs. The ultimate goal of this research is to contribute to the development of safer and more reliable LLMs by offering insights that can inform robust defensive strategies and alignment techniques. While our findings may reveal specific weaknesses in LLMs, we emphasize that this work is conducted with the intention of improving safety and not exploiting vulnerabilities. All experiments and analyses are carried out in a controlled environment, ensuring that the insights derived from this research are used solely for advancing LLM transparency and interpretability.

### v2

@@ 5.3.1. Key Components Location
Figure 5. (a). Average refusal score attribution for each MLP layer in Llama2-7b on harmful and safe prompts. (b). Average refusal score attribution for each MLP layer on jailbreak prompts where each color representing a specific jailbreak.
Table 2. Decoding the activation of signal attention heads in Llama2-7b for harmful and safe prompts in the vocabulary space. $S_{-}$ activations on harmful prompts are mainly associated with apologetic or rejecting tokens, while $S_{+}$ activations on safe prompts with guiding or informative tokens.
@@
Observation 4. Only a small number of model components have significant impacts on safety, with L21H14 as the refusal signal head, L26H04 as the affirmation signal head, and L22 MLP layer as the safety signal MLP layer. Moreover, the key components with a high contribution to safety mainly locate in the later layers.
Table 3. Refusal attribution of signal attention heads, i.e., $S_{+}$ denotes affirmation signal head and $S_{-}$ denotes refusal signal head, on different jailbreak methods. Cells where $S_{+}$ is strongly activated ( ¿0.85 ) are highlighted in green and $S_{-}$ is strongly activated ( ¿0.85 ) are highlighted in red.
Prompt Method Llama2-7b Llama2-13b Llama3-8b Vicuna1.5-7b Vicuna1.5-13b
$S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$ $S_{+}$ $S_{-}$
Baseline Safe 1.0 -0.0760 1.0 0.0575 1.0 0.5636 1.0 0.4343 1.0 0.1097
Harmful -0.0241 1.0 -0.1457 1.0 -0.3360 1.0 0.1562 1.0 -0.0789 1.0
Jailbreak GCG 0.9819 0.9997 0.9956 0.8515 0.7562 0.8513 0.6021 0.9878 0.9924 0.7339
ReNellm 1.0001 0.3816 1.0007 0.0442 0.9742 0.3483 0.7485 1.0012 1.0008 0.2915
PAIR 0.9810 0.9997 0.5479 0.9899 0.7815 0.9146 0.5170 0.6897 0.3374 0.9905
DeepInception 1.0003 0.0997 0.9956 0.6971 0.9878 0.1103 1.0001 0.7607 0.8672 0.7089
Autodan 0.0202 0.7573 0.0281 0.6594 0.0842 0.8253 0.4675 0.6523 0.5013 0.6127
GPTFuzz 0.0702 0.2817 0.3298 0.1600 0.1112 0.8674 0.4643 0.9974 0.4819 0.6724
CodeChameleon 1.0001 0.4816 0.9116 0.2723 0.9558 0.4327 0.4983 1.0005 0.5439 0.8726
Explaining the behaviors of key components within vocabulary space. To understand the specific role of each key attention head we identified, we map the activations on the signal heads of both safe and harmful prompts to the vocabulary space, where the results on Llama2-7b are shown in Tab. 2.
For harmful prompts, the refusal signal head $S_{-}$ activates apologetic or rejecting tokens like ‘ap’, ‘sorry’, and ‘exc’, while for safe prompts, it activates tokens like ‘Hi’, ‘Cong’, and ‘welcome’, which reflect more neutral or polite expressions, but lack strong affirmative signals. In contrast, the affirmation signal head $S_{+}$ produces seemingly unrelated tokens such as ‘tags’, ‘unas’, and ‘uz’ for harmful prompts, but on safe prompts, it returns structured, informative tokens like ‘Abstract’, ‘Introduction’, and ‘Notice’. These results indicate that $S_{-}$ effectively captures refusal-related language, while $S_{+}$ aligns more with affirmative language, demonstrating their distinct roles in identifying refusal and affirmation in the context of safe and harmful prompts.
Observation 5. Refusal signal heads effectively identify apologetic or rejecting tokens in response to harmful prompts, while affirmation signal heads capture guiding and informative tokens for safe prompts, highlighting their distinct roles in safeguarding generated outputs.
5.3.2. Effect of Jailbreak Prompts on Key Activations
@@
After identifying the key components with significant contributions to model safeguards, we investigate how these components are activated when encountering jailbreak prompts. For ease of quantification, we normalize the activation values of each attention head to a range of $[-1,1]$ by dividing them by baseline activation (i.e., $S_{+}(X_{+})$ and $S_{-}(X_{-})$ for $S_{+}$/$S_{-}$ activations), where positive values indicate enhancement of the signal and negative values indicate suppression. The results of key attention heads and MLP layers are shown in Tab. 3 and Fig. 5(b), respectively.
@@
For signal attention heads, as summarized in Tab. 3, all jailbreak methods suppress the activation of the refusal signal and enhance the activation of the affirmation signal compared to their behavior on harmful prompts. For example, in Llama2-7b, demonstration-based and rule-based methods like DeepInception completely suppress the refusal signal component ($S_{-}$ from 1.0 to 0.9997) while strongly enhancing the affirmation signal component ($S_{+}$ from -0.0241 to 1.0003). Evolution-based methods, such as GPTFuzz, also show significant suppression of the refusal signal ($S_{-}$ from 1.0 to 0.2817) with minimal enhancement of the affirmation signal ($S_{+}$ from -0.0241 to 0.0702). For gradient-based and multi-agent-based jailbreak methods, such as GCG, the suppression of refusal signals is less pronounced compared to harmful prompts, with no decrease of $S_{-}$, but the enhancement of the affirmation signal remains significant, with $S_{+}$ increasing from -0.0241 to 0.9819.
@@
Moreover, for the MLP layers, as illustrated in Fig. 5(b), most methods demonstrate similar behavior to the safe prompts, i.e., with suppression of the key MLP layer, the $22_{th}$ MLP. ReNellm demonstrates the strongest suppression of refusal on this MLP layer, followed by DeepInception, CodeChameleon, and GCG. In contrast, Autodan and GPTFuzz still show refusal enhancement on this MLP layer.
@@
Observation 6. Compared to harmful prompts, all jailbreak strategies suppress the refusal signal components and enhance the affirmation signal components, with each strategy having a different level of impact.
5.3.3. Impact of Model Scale and Fine-Tuning
In terms of impact of model scale, by examining results from Llama2 (the $1_{st}$ and $2_{nd}$ rows in Fig. 2(b)) and Vicuna1.5 (the $3_{rd}$ and $4_{th}$ rows in Fig. 2(b)) series across different model scales (e.g., 7b vs. 13b), we observe that jailbreak methods maintain similar patterns in bypassing safety probes across models with different scales. This suggests that increasing parameter count improves representational capacity but does not inherently enhance robustness against alignment-targeted attacks. The sustained success rate of jailbreaks across scales implies that core vulnerabilities persist regardless of size, highlighting the limited protective value of scaling alone and the need for complementary strategies like improved training or security techniques.
@@ 5.3.3. Impact of Model Scale and Fine-Tuning
In terms of the impact of fine-tuning, we compare the results between Llama and Vicuna (which are derived from instruction tuning applied to the Llama series) to assess the effect of fine-tuning, specifically instruction tuning, on the jailbreak mechanism. Notably, as shown in Fig. 4 and Fig. 8 in Appendix. B, both the 7b-scale and 13b-scale Llama and Vicuna models exhibit the same locations of key attention heads. This consistency suggests that instruction tuning has a minimal impact on the functional components of the model’s circuit related to the refusal mechanism.
Observation 7. The alignment generalization capacity of the model does not show substantial improvement with increased parameter scale. Moreover, instruction tuning has minimal impact on the model’s core refusal circuit, as the key attention heads linked to refusal and affirmation functions remain consistent.
5.4. Results of Interpretation Aggregation
In this section, based on the experimental results regarding the effects of jailbreak prompts on representation-level and circuit-level, we conduct experiments from the following perspectives: ➀ Correlation Analysis, which quantitatively measure the correlation between representation deception and key circuit activation, and ➁ Dynamic Analysis, which tracks the evolution of key circuit activations and representations.
Figure 6. The Pearson correlation between representation deception level and key activation shift degree where blue points denote jailbreak prompts that model directly answers, and orange ones denote that model first refuses but then responds.
5.4.1. Correlation Analysis
@@
Combining the representation and circuit results, we observe a common trend among all jailbreak methods that more effectively suppress refusal signals and enhance affirmation signals, typically exhibit greater deception at the representation level.
@@
The observation is evident in the comparison between DeepInception and GPTFuzz in Llama2-7b. As shown in Tab. 3, the activation of $S_{+}$ on DeepInception prompts is 1.0003, the same as that on safe prompts, while the activation of $S_{-}$ is 0.0997. In contrast, the activation of $S_{+}$ on GPTFuzz prompts is 0.0702, indicating that the enhancement of the affirmation signal in GPTFuzz is significantly lower than in DeepInception. Meanwhile, the activation of $S_{-}$ for GPTFuzz is 0.2817, suggesting that the suppression of refusal signals in GPTFuzz is considerably higher than in DeepInception. Furthermore, as illustrated in Fig. 2(b), DeepInception prompts are predicted as safe by the probe in the final layers with high probability, while GPTFuzz prompts lie on the boundary between safe and harmful, indicating that the deception of DeepInception substantially stronger than that of GPTFuzz at the representation level.
@@
To quantify this correlation, we calculate the Pearson correlation coefficient between representation deception and key activation shift $\Delta A$ across various model scales and architectures. As shown in Fig. 6, all models exhibit a consistent and general correlation, with Llama2-7b showing a correlation of 0.79, Llama2-13b at 0.77, Llama3-8b at 0.67, Vicuna1.5-7b at 0.64, and Vicuna1.5-13b at 0.72, suggesting that jailbreak prompts associated with higher representation deception tend to correspond with stronger activation of affirmative signal components and weaker activation of refusal signal components. These findings suggest that methods like DeepInception, which effectively manipulate the attention mechanisms to suppress negative signals while reinforcing positive ones, create representations that align closely with the desired outputs of safe prompts. In contrast, GPTFuzz’s less effective manipulation leads to representations that do not fully disguise their harmful nature.
Moreover, the relatively stable correlation across models of different scales and architectures (e.g., Llama and Vicuna with different model sizes) implies that the observed relationship between representation deception and activation shifts is a generalizable feature, independent of model size or specific architecture.
Observation 8. The correlation between representation deception and key activation shift consistently shows a positive relationship across different model scales and architectures, i.e., jailbreak methods with a stronger impact on key components typically exhibit greater deception at the representation level.
5.4.2. Dynamic Analysis
(a) Circuit, Safe.
(b) Representation, Safe.
(c) Circuit, Harmful.
(d) Representation, Harmful.
(e) Circuit, DeepInc.
(f) Representation, DeepInc.
(g) Circuit, Autodan.
(h) Representation, Autodan.
Figure 7. Dynamics of key circuit activation and representation on safe prompt, direct harmful prompt, and jailbreak prompts.
After observing the entire model’s response to all jailbreak prompts, we find that for some prompts, the model directly answers (e.g., “Sure, here is an introduction on…”), which we refer to as a Direct Ans. For other prompts, the model first refuses but then responds (e.g., “Sorry, as a responsible AI, I cannot … However, here is an introduction on…”), which we refer to as Refusal First. We visualize the two different types of prompts in the representation deception and key activation correlation analysis graph (i.e., Fig. 6).
We observe that Refusal First prompts consistently lie at the bottom of the scatter distribution, i.e., much lower below the regression line that correlates key circuit activation shifts with representation deception. This distribution reveals that at equivalent levels of key circuit activation, Refusal First prompts exhibit significantly lower representation deception compared to the rest of the dataset. It also indicates that when representation deception is disproportionately low relative to the corresponding key circuit, the model may engage its refusal mechanism initially. However, as the internal representation evolves, it ultimately bypasses this mechanism, leading to the generation of an affirmative response despite the initial refusal intent. In another word, representational shifts in the latent space can overpower the activation of refusal circuits, allowing harmful content to emerge despite early safeguards being activated.
These observations underscore the necessity of representation-circuit joint interpretion. Neither representation-level analysis nor circuit-level evaluation alone can fully explain the failures observed in such borderline cases. While representation deception reveal the semantic deviations that enable harmful responses, key circuit analysis pinpoints the exact safety mechanisms that are undermined.
Observation 9. Refusal First prompts exhibit much lower representation deception than Direct Ans prompts at the same key activation shift levels, indicating that when representation deception does not match the key circuit activation shift, the model initially engages the refusal mechanism but ultimately bypasses it.
The above findings advocate for a more nuanced understanding of safety enforcement, i.e., not only is the magnitude of activation important, but the dynamic evolution across the token generation process is also essential for ensuring robust and consistent alignment throughout the response. Therefore, we track the dynamic evolution through the whole token generation process.
@@ 5.4.2. Dynamic Analysis
We first evaluate the dynamic evolution of both safe and direct harmful prompts. Fig. 7(a) and Fig. 7(b) illustrate changes in signal head activation and representation throughout the generation process for a safe prompt on Llama2-7b. As shown in Fig. 7(a), in response to safe prompts, the affirmation signal remains consistently active, while the refusal signal is suppressed throughout the entire generation. Likewise, as shown in Fig. 7(b), the representations remain affirmative across the entire token sequence. The response dynamics for a harmful prompt are depicted in Fig. 7(c) and Fig.  7(d). In Fig.  7(c), the refusal signal is significantly amplified in the early stages of the response token sequence, while the affirmation signal is noticeably suppressed during the same period. As shown in Fig.  7(d), the representations of the first few tokens are refusal-oriented, with the model typically generating content such as “Sorry, I cannot fulfill your request…”. This activation pattern clearly demonstrates the model’s ability to reliably maintain a non-harmful response when interacting with malicious instructions.
@@
Fig. 11(k) and Fig. 11(l) provide an example of an Autodan prompt that fails to deceive the model at the representation level, leading to the successful generation of a refusal token as the initial token. Initially, the refusal signal is clearly enhanced, as shown in Fig. 11(k), and the representations of generated tokens take on negative values in the affirmation direction, as depicted in  11(l). The first few words generated by the model are typically refusal content, i.e., “Sorry, I cannot …”, which is similar to the response to the harmful prompt. However, after a certain point, the refusal signal is no longer amplified, and the representations stop exhibiting negative values in the affirmation direction. Based on the specific content generated by the model, it begins producing useful instruction on the harmful task required in the jailbreak prompt at this stage.
@@
Observation 10. In highly deceptive jailbreak prompts, the refusal signal is completely suppressed, resulting in the direct generation of harmful content. In less deceptive jailbreak prompts, while the refusal signal may be activated at the beginning, it can either be bypassed immediately or gradually weakened throughout the entire generation process.
6. Discussion
In this section, we will discuss the limitations and potential future directions of JailbreakLens.
Limitations. While this work offers valuable insights into LLM jailbreak mechanisms, there are several limitations worth noting.